OpenStudy (123ab456c):

In a jar of coins, the ratio of nickels to dimes is 2:7. If the jar contains 350 nickels, how many dimes are there? A) 175 B) 350 C) 700 D) 1225

rishavraj (rishavraj):

say no of nickles = 2x and tht of dimes = 7x so 2x = 350 get x and the no of dimes

rishavraj (rishavraj):

so like 2x = 350 divide by 2 on both sides so x = 175 ... and no of dimes = 7*x = 7*175 = ???

OpenStudy (mathmale):

Ratios: nickels 2 nickels to dimes: ---------- = --- = dimes 7 That means the very smallest number of nickels you could have would be 2 and the very smallest number of dimes 7. Add 2 to 7. Result: Fraction of the 350 total coins that are nickels is 2/9; fraction that are dimes is 7/9. Multiply 350 by 2/9. # of nickels is ... ?